# A New Perspective to Fish Trajectory Imputation: A Methodology for Spatiotemporal Modeling of Acoustically Tagged Fish Data

### v1

@@ 6 Discussion
In summary, this paper focuses on the imputation of cobia movement trajectories while accounting for physical limitations due to the tracking technology (acoustic tagging), accompanied by visual validation. The challenges addressed in this paper are the structure of the data and the methods used for data collection temporally. The data are very sparse, and the actual fish location is unknown even when presence is detected. In fact, these are the complexities of tracking the fish acoustically, which lead to a lack of applicability of the common models for animal movement analysis. In our paper, we developed a new strategy to handle these challenges.
@@ Acknowledgement
We gratefully acknowledge the financial support of the Virginia Sea Grant through a graduate fellowship awarded to M.A. which made this research possible. We also extend our appreciation to Dr. Daniel P. Crear (Inter-American Tropical Tuna Commission) and Dr. Kevin Weng (Virginia Institute of Marine Science) for providing the necessary resources and their guidance during this work.
@@ Appendix A Second-order Taylor Expansion for Z
$\approx(\sin\theta)(\theta^{*}-\theta)-\cos\theta\left(1-\frac{(\theta^{*}-\theta)^{2}}{2}\right)$
$\mathbb{E}(Z^{(1)})$ $\approx\cos\theta\left(1-\frac{\sigma_{\psi}^{2}}{2}\right)$
$\text{Var}(Z^{(1)})$ $\approx(\sin\theta)^{2}\sigma^{2}_{\psi}+\frac{1}{2}(\cos\theta)^{2}\sigma^{4}_{\psi}$
@@
$\approx\cos\theta(\theta^{*}-\theta)-(\sin\theta)\left(1-\frac{(\theta^{*}-\theta)^{2}}{2}\right)$
$\mathbb{E}(Z^{(2)})$ $\approx(\sin\theta)\left(1-\frac{\sigma_{\psi}^{2}}{2}\right)$
$\text{Var}(Z^{(2)})$ $\approx(\cos\theta)^{2}\sigma^{2}_{\psi}+\frac{1}{2}(\sin\theta)^{2}\sigma^{4}_{\psi}$
@@
$\text{Cov}(Z^{(1)},Z^{(2)})\approx-\mathrm{S}\mathrm{C}\sigma^{2}_{\psi}+\frac{1}{2}\mathrm{S}\mathrm{C}\sigma^{4}_{\psi}$
@@ Appendix C High resolution heat map
Download Figure 2 in high resolution HTML format at https://go.wm.edu/4YYgLP.

### v2

@@ 6 Discussion
In summary, this paper focuses on the imputation of cobia movement trajectories while accounting for physical limitations due to the tracking technology (acoustic tagging), accompanied by visual validation. Our imputation approach relies on simulating from bespoke probability distributions tailored to the data collection technology and animal species of interest. The challenges addressed in this paper are the structure of the data and the methods used for data collection temporally. The data are very sparse, and the actual fish location is unknown even when presence is detected. In fact, these are the complexities of tracking the fish acoustically, which lead to a lack of applicability of the common models for animal movement analysis. In our paper, we developed a new strategy to handle these challenges.
@@ Acknowledgement
We gratefully acknowledge the financial support of the Virginia Sea Grant through a graduate fellowship awarded to M.A. which made this research possible. We also extend our appreciation to Dr. Daniel P. Crear (Inter-American Tropical Tuna Commission) and Dr. Kevin Weng (Virginia Institute of Marine Science) for providing the necessary resources and their guidance during this work. We thank the handling editor for helpful comments that improved our paper.
@@ Appendix A Second-order Taylor Expansion for Z
$\approx(\sin\theta)(\theta^{*}-\theta)-\cos\theta\left(1-\frac{(\theta^{*}-\theta)^{2}}{2}\right),$
$\mathbb{E}(Z^{(1)})$ $\approx\cos\theta\left(1-\frac{\sigma_{\psi}^{2}}{2}\right),$
$\text{Var}(Z^{(1)})$ $\approx(\sin\theta)^{2}\sigma^{2}_{\psi}+\frac{1}{2}(\cos\theta)^{2}\sigma^{4}_{\psi}\,,$
@@
$\approx\cos\theta(\theta^{*}-\theta)-(\sin\theta)\left(1-\frac{(\theta^{*}-\theta)^{2}}{2}\right),$
$\mathbb{E}(Z^{(2)})$ $\approx(\sin\theta)\left(1-\frac{\sigma_{\psi}^{2}}{2}\right),$
$\text{Var}(Z^{(2)})$ $\approx(\cos\theta)^{2}\sigma^{2}_{\psi}+\frac{1}{2}(\sin\theta)^{2}\sigma^{4}_{\psi}\,,$
@@
$\text{Cov}(Z^{(1)},Z^{(2)})\approx-\mathrm{S}\mathrm{C}\sigma^{2}_{\psi}+\frac{1}{2}\mathrm{S}\mathrm{C}\sigma^{4}_{\psi}\,,$
@@ Appendix B Joint Distribution of Stochastic Terms for $T=4$
$\mathcal{P}\left[\begin{matrix}X_{1}\end{matrix}\right]\mathcal{P}\left[\begin{matrix}X_{4}\end{matrix}\right].$
@@ Appendix C High resolution heat map
Download Figure 2 in high resolution HTML format at https://go.wm.edu/NYq32h.